# Continuum Attention for Neural Operators

### v1

@@ 2.1.2 Sequences over $D\subset\mathbbm{R}^{d}$
We note that the preceding two formulations of the attention mechanism view it as a transformation defined as an expectation applied to the input sequence. The expectation is with respect to a probability measure supported on an index of the input sequence; furthermore the expectation itself depends on the input sequence to which it is applied, and is parameterized by the input sequence index, resulting in a new output sequence of the same length as the input sequence. The formulation using an expectation results from the softmax part of the definition of the attention mechanism; the query and key linear operations define the expectation, and the value linear operation lifts the output sequence to take values in a (possibly) different Euclidean space than the input sequence. These components may be readily extended to work with (generalized) sequences defined over open subsets of $\mathbbm{R}^{d}$. (These would often be referred to as functions, but we call them generalized sequences to emphasize similarity with the discrete case.)
@@
This result connects our continuum formulation of self-attention with the standard definition, using Monte Carlo approximation of the relevant integrals; a similar result could be proved using finite difference approximation on, for example, a uniform grid.
@@ 2.2.2 Sequences over $D\subset\mathbbm{R}^{d},E\subset\mathbbm{R}^{e}$
Proof  The proof of this result is developed in Appendix A.2.
@@ 3 Continuum Patched Attention
Throughout this section, we let $D\subset\mathbbm{R}^{d}$ be a bounded open set and let $D\coloneqq D_{1}\cup\ldots\cup D_{P}$ be a uniform partition of the space $D$ so that $D_{j}\cong D^{\prime}$ are congruent for all $j\in\llbracket 1,P\rrbracket$ for some $D^{\prime}\subset D$, where $P\in\mathbb{N}$ represents the number of patches. We consider a function defined on the full domain $D$, denoted as $u\in\mathcal{U}\bigl{(}D;\mathbbm{R}^{d_{u}}\bigr{)}$. We define a mapping
@@
$\widetilde{u}(p)={\left.\kern-1.2ptu\mathchoice{\vphantom{\big{|}}}{}{}{}\right|_{D_{p}}},$ (6)
@@ 3.1 Patched Self-Attention
The self-attention operator $\mathsf{A}_{\mathsf{P}}$ maps the function taking a patch index to a $\mathbbm{R}^{d_{u}}-$valued function over $D^{\prime}$, $\widetilde{u}$, into a function taking a patch index to a $\mathbbm{R}^{d_{V}}-$valued function over $D^{\prime}$, $\mathsf{A}_{\mathsf{P}}(\widetilde{u})$, as follows:
@@
Letting $N$ be the total number of discretization points in $D$ and $P$ the number of patches, in the discrete setting each function patch $\widetilde{u}(p):D^{\prime}\to\mathbbm{R}^{d_{u}}$ may be represented as a $\mathbbm{R}^{N/P\times d_{u}}$ matrix. The flattening procedure of each patch in ViT involves reshaping this matrix into a $\mathbbm{R}^{N/P\cdot d_{u}}$ vector. The full image is thus represented as a sequence of $P$ vectors of dimension $N/P\cdot d_{u}$. Each patch vector is then linearly lifted to an embedding space of dimension $d_{\textrm{model}}$. The embedded image, to which attention is applied, may hence be viewed as a matrix $\widetilde{u}\in\mathbbm{R}^{P\times d_{\textrm{model}}}$. Therefore, ViT may be cast in the setting of the discrete interpretation of the self-attention operator outlined in Subsection 2.1.1, where the domain on which attention is applied is $D^{P}$. A key observation to the ensuing discussion is that the application of a linear transformation to a “flattened” patch in ViT breaks the mesh-invariance of the architecture; however, this operation may be viewed as a nonlocal transformation on each patch. We leverage this insight in Section 4 to design discretization invariant transformer neural operators that employ patching.
@@ 3.2 Patched Cross-Attention
$\widetilde{v}(o)={\left.\kern-1.2ptv\mathchoice{\vphantom{\big{|}}}{}{}{}\right|_{E_{o}}},$ (8)
@@
The cross-attention operator $\mathsf{C}$ maps the operator taking a patch index to a $\mathbbm{R}^{d_{u}}-$valued function over $D^{\prime}$, $\widetilde{u}$, and the operator taking a patch index to a $\mathbbm{R}^{d_{v}}-$valued function over $E^{\prime}$, $\widetilde{v}$, into an operator taking a patch index to a $\mathbbm{R}^{d_{V}}-$valued function over $D^{\prime}$, $\mathsf{C}_{\mathsf{P}}(\widetilde{u},\widetilde{v})$, as follows:
@@ 4 Transformer Neural Operators
In this section we turn our attention to describing how to devise transformer-based neural operators using the continuum attention operator $\mathsf{A}$ as defined in the function space setting in Section 2, and the continuum patched-attention operator $\mathsf{A}_{\mathsf{P}}$ acting on function space from Section 3. We make use of the operator frameworks developed in Sections 2 and 3 to formulate transformer neural operator architectures acting on function space. We start in Subsection 4.1, setting-up the framework. In Subsection 4.2 we formulate an analogue of the transformer from Vaswani et al. (2017) as an operator mapping an input function space $\mathcal{U}\bigl{(}D;\mathbbm{R}^{d_{u}}\bigr{)}$ to the solution function space $\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}$. The resulting scheme is invariant to the discretization of the input function and allows for zero-shot generalization to irregular, non-uniform meshes. Motivated by the need to develop new efficient architectures to employ attention on longer sequences, or in our case higher resolution discretizations of functions, in Subsection 4.3 we describe a generalization of the ViT methodology to the function space setting, inspired by the work in Dosovitskiy et al. (2021). In particular we generalize the procedure of lifting flattened patches to the embedding dimension, as done in ViT, to the function space setting. This is achieved by lifting patches of functions to an embedding dimension using a nonlocal linear operator, namely, an integral operator. Unlike ViT, the resulting neural operator architecture is discretization invariant and allows for zero-shot generalization to different resolutions. In Subsection 4.4 we describe a different patching-based transformer neural operator acting on function space. The architecture is discretization invariant and is based on $\mathsf{Q},\mathsf{K},\mathsf{V}$, learnable parameters in the attention mechanism, being defined as integral operators acting themselves on function space.
@@ 4.1 Set-Up
for any $v\in\mathcal{V}$. Here the notation $\mathsf{A}_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ indicates the ambiguity of the definition of self-attention. We note that we will make explicit which definition of self-attention operator from the previous sections we employ for each architecture; we will also make clear the definitions of the parameters $\theta_{h}\in\Theta$ defining each attention operator. The operators $\mathsf{W}_{1},\mathsf{W}_{2}:\mathcal{V}\to\mathcal{V}$ are pointwise linear operators.  [^4] We note that for all experiments in Section 6 we set $\mathsf{W}_{1},\mathsf{W}_{2}$ to be the identity operators. Next, the operator $\mathsf{F}_{\textrm{LayerNorm}}:\mathcal{V}\to\mathcal{V}$ defines the layer normalization and the operator $\mathsf{F}_{\textrm{NN}}:\mathcal{V}\to\mathcal{V}$ defines a feed-forward neural network layer. Both operators are applied pointwise to the input; further details are provided in the context of each architecture.
@@
Assuming ${\kappa}_{\theta}(x,y)={\kappa}_{\theta}(x-y)$ then (12) may be viewed as a convolution operator, which makes it possible to use the Fast Fourier Transform (FFT) to compute (12) and to parametrize ${w}_{\theta}$ in Fourier space. Following Li et al. (2021), this insight leads to the following formulation of the operator given in Definition 19.
@@
For uniform discretizations $n_{1}\times\ldots\times n_{d}=N$ of the space $D$, the Fourier transform can be replaced with the Fast Fourier Transform (FFT). Indeed, for the Fourier transform we have $\mathcal{F}u:\mathbbm{R}^{d}\to\mathbbm{C}^{r}$ and $R_{\mathsf{W}}(\theta):\mathbbm{R}^{d}\to\mathbbm{C}^{r^{\prime}\times r}$. On the other hand, for the FFT we have $\mathcal{F}u:\mathbbm{Z}^{d}\to\mathbbm{C}^{r}$ and $R_{\mathsf{W}}(\theta):\mathbbm{Z}^{d}\to\mathbbm{C}^{r^{\prime}\times r}$. In practice, it is possible to select a finite-dimensional parametrization by choosing a finite set of wave numbers
$\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}:=\bigl{|}\bigl{\{}k\in\mathbbm{Z}^{d}:|k_{j}|\leq k_{\textrm{max},j},\,\text{for}\,j=1,\ldots,d\bigr{\}}\bigr{|}.$
We can therefore parametrize $R_{\mathsf{W}}$ by a complex $(2k_{\text{max},1}\times\ldots\times 2k_{\text{max},d}\times r^{\prime}\times r)$-tensor.
@@ 4.2 Transformer Neural Operator
Next, we define the details of $\mathsf{E}_{L}:\mathcal{V}\to\mathcal{V}$ in Equation 10 by specifying a definition of $\mathsf{A}_{\textrm{MultiHead}}$, $\mathsf{W}_{1},\mathsf{W}_{2}$, $\mathsf{F}_{\textrm{LayerNorm}}$, and $\mathsf{F}_{\textrm{NN}}$ for the transformer neural operator setting. The operator $\mathsf{A}_{\textrm{MultiHead}}:\mathcal{V}\to\mathcal{V}$ is the multi-head attention operator from (11) based on the continuum self-attention from Definition 5. Letting $H\in\mathbbm{N}$ denote the number of attention heads, the multi-head attention operator is parametrized by the learnable linear transformations $Q_{h}\in\mathbbm{R}^{d_{K}\times d_{\textrm{model}}},K_{h}\in\mathbbm{R}^{d_{K}\times d_{\textrm{model}}},V_{h}\in\mathbbm{R}^{d_{K}\times d_{\textrm{model}}}$, for $h=1,\ldots,H$, where for implementation purposes $d_{K}$ is chosen as $d_{K}\coloneqq d_{\textrm{model}}/H$.  [^5] We note that $d_{\textrm{model}}$ and $H$ should be chosen so that $d_{\textrm{model}}$ is a multiple of $H$. The multi-head attention operator is thus defined by the application of a linear transformation $W_{\textrm{MultiHead}}\in\mathbbm{R}^{d_{\textrm{model}}\times H\cdot d_{K}}$ to the concatenation of the outputs of $H\in\mathbbm{N}$ self-attention operators. The operators $\mathsf{W}_{1},\mathsf{W}_{2}:\mathcal{V}\to\mathcal{V}$ are pointwise linear operators and hence admit finite-dimensional representations $W_{1},W_{2}\in\mathbbm{R}^{d_{\textrm{model}}\times d_{\textrm{model}}}$, so that they define a map
@@
for $k=1,\dots,d_{\textrm{model}}$, any $v\in\mathcal{V}$, and any $x\in D\subset\mathbbm{R}^{d}$, where the notation $(\,\cdot\,)_{k}$ is used to denote the $k$’th entry of the vector. In equation (16), ϵ+italic-$\epsilon\in\mathbbm{R}^{+}$ is a fixed parameter, $\gamma_{k},\beta_{k}\in\mathbbm{R}$ are learnable parameters and $m,\sigma$ are defined as
$m(y)$ $\coloneqq\frac{1}{d_{\textrm{model}}}\cdot|y|_{1},$ (17)
@@ 4.3 Vision Transformer Neural Operator
Let $D\subset\mathbbm{R}^{d}$ be a bounded open set. Consider $u\in\mathcal{U}(D;\mathbbm{R}^{d_{u}})$ and let $D\coloneqq D_{1}\cup\ldots\cup D_{P}$ be a uniform partition of the space $D$ so that $D_{j}\cong D^{\prime}$ for all $j\in\llbracket 1,P\rrbracket$ for some $D^{\prime}\subset D$, where $P\in\mathbb{N}$ represents the number of patches. We consider model inputs $u\in\mathcal{U}(D;\mathbbm{R}^{d_{u}})$, model embedding space $\mathcal{V}:=\left\{v\mid v:\llbracket 1,P\rrbracket\to\mathcal{U}\bigl{(}D^{\prime};\mathbbm{R}^{d_{\textrm{model}}}\bigr{)}\right\}$ which is a space of functions acting on patch indices, and model output space $\mathcal{Z}(D;\mathbbm{R}^{d_{z}})$, with domain $D\subset\mathbbm{R}^{d}$.
@@
for $k=1,\dots,d_{\textrm{model}}$, any $p\in\llbracket 1,P\rrbracket$ and any $x\in\mathbbm{R}^{d}\subset D$, where we use the notation $(\,\cdot\,)_{k}$ to denote the $k$’th entry of the vector. In equation (24), ϵ+italic-$\epsilon\in\mathbbm{R}^{+}$ is a fixed parameter, $\gamma_{k},\beta_{k}\in\mathbbm{R}$ are learnable parameters and $m,\sigma$ are defined as in (17). The operator $\mathsf{F}_{\textrm{NN}}:\mathcal{V}\to\mathcal{V}$ is defined by
@@
We have found that adding a convolution in residual form helps to ameliorate the discontinuities arising when the patching architecture is used; these are visible in error plots and can also interfere with self-composition of maps learned as solution operators of time-dependent PDEs. To this end define the linear integral operator $\mathsf{W}_{\textrm{S}}\in\mathcal{L}\Bigl{(}\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}\,;\,\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}\Bigr{)}$, of the form described in Definition 20 and the subsequent discussion, and add a last layer to the output of the architecture as defined thus far:
${z}\mapsto z+\mathsf{W}_{\rm{S}}z\in\mathcal{Z}(D;\mathbbm{R}^{d_{z}}),$ (29)

### v2

@@ 2.1.2 Sequences over $D\subset\mathbbm{R}^{d}$
We note that the preceding two formulations of the attention mechanism view it as a transformation defined as an expectation applied to the input sequence. The expectation is with respect to a probability measure supported on an index of the input sequence. Furthermore, the expectation itself depends on the input sequence to which it is applied, and is parameterized by the input sequence index, resulting in a new output sequence of the same length as the input sequence. The formulation using an expectation results from the softmax part of the definition of the attention mechanism; the query and key linear operations define the expectation, and the value linear operation lifts the output sequence to take values in a (possibly) different Euclidean space than the input sequence. These components may be readily extended to work with (generalized) sequences defined over open subsets of $\mathbbm{R}^{d}$. (These would often be referred to as functions, but we call them generalized sequences to emphasize similarity with the discrete case.)
@@
This result connects our continuum formulation of self-attention with the standard definition, using Monte Carlo approximation of the relevant integrals; a similar result could be proved using finite difference approximation on, for example, a uniform grid. We note that it is not possible to obtain a uniform convergence rate for all $u\in B$ in the norm of $C(\bar{D};\mathbb{R}^{d_{V}})$ as the Riemann integral may exhibit arbitrarily slow convergence depending on the regularity of the integrand.
@@ 2.2.2 Sequences over $D\subset\mathbbm{R}^{d},E\subset\mathbbm{R}^{e}$
Proof  The proof of this result is developed in Appendix A.2, which contains straightforward generaliations of the techniques in Appendix A.1.
@@ 3 Continuum Patched Attention
Throughout this section, we let $D\subset\mathbbm{R}^{d}$ be a bounded open set and let $D\coloneqq D_{1}\cup\cdots\cup D_{P}$ be a uniform partition of the space $D$ so that $D_{j}\cong D^{\prime}$ are congruent for all $j\in\llbracket 1,P\rrbracket$ for some $D^{\prime}\subset D$, where $P\in\mathbb{N}$ represents the number of patches. We note that we require a uniform partition, namely a partition of subsets of equivalent finite volume, in order to define operators that can be applied to functions defined on each of these subsets; however, for practical application we may have a different number of discretization points within each subset. We consider a function defined on the full domain $D$, denoted as $u\in\mathcal{U}\bigl{(}D;\mathbbm{R}^{d_{u}}\bigr{)}$. We define a mapping
@@
[…]
@@ 3.1 Patched Self-Attention
[…]
@@
Letting $N$ be the total number of discretization points in $D$ and $P$ the number of patches, in the discrete setting each function patch $\widetilde{u}(p):D^{\prime}\to\mathbbm{R}^{d_{u}}$ may be represented as an $\mathbbm{R}^{N/P\times d_{u}}$ matrix. The flattening procedure of each patch in ViT involves reshaping this matrix into an $\mathbbm{R}^{N/P\cdot d_{u}}$ vector. The full image is thus represented as a sequence of $P$ vectors of dimension $N/P\cdot d_{u}$. Each patch vector is then linearly lifted to an embedding space of dimension $d_{\textrm{model}}$. The embedded image, to which attention is applied, may hence be viewed as a matrix $\widetilde{u}\in\mathbbm{R}^{P\times d_{\textrm{model}}}$. Therefore, ViT may be cast in the setting of the discrete interpretation of the self-attention operator outlined in Subsection 2.1.1, where the domain on which attention is applied is $D^{P}$. A key observation to the ensuing discussion is that the application of a linear transformation to a “flattened” patch in ViT breaks the mesh-invariance of the architecture; however, this operation may be viewed as a nonlocal transformation on each patch. We leverage this insight in Section 4 to design discretization invariant transformer neural operators that employ patching. We further note that in practice, patching in ViT is achieved via strided convolutions, with stride equivalent to the kernel size; however, because the kernel size determines the number of parameters, such a procedure also breaks discretization invariance as it leads to architectures with parameters dependent on resolution of the input.
@@ 3.2 Patched Cross-Attention
[…]
@@
[…]
@@ 4 Transformer Neural Operators
In this section we turn our focus to describing how to devise transformer-based neural operators using the continuum attention operator $\mathsf{A}$ as defined in the function space setting in Section 2, and the continuum patched-attention operator $\mathsf{A}_{\mathsf{P}}$ acting on function space from Section 3. We make use of the operator frameworks developed in Sections 2 and 3 to formulate transformer neural operator architectures acting on function space. We start in Subsection 4.1, setting-up the framework. In Subsection 4.2 we formulate an analogue of the transformer from Vaswani et al. (2017) as an operator mapping an input function space $\mathcal{U}\bigl{(}D;\mathbbm{R}^{d_{u}}\bigr{)}$ to the solution function space $\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}$. The resulting scheme, which we define as transformer neural operator (TNO) is invariant to the discretization of the input function and allows for zero-shot generalization to irregular, non-uniform meshes. However, this architecture exhibits quadratic complexity with respect to the number of discretization points of the input functions. Hence, motivated by the need to develop new efficient architectures to employ attention on longer sequences, or in our case higher resolution discretizations of functions, in Subsection 4.3 we describe a generalization of the ViT methodology to the function space setting, inspired by the work in Dosovitskiy et al. (2021). In particular we generalize the procedure of lifting flattened patches to the embedding dimension, as done in ViT, to the function space setting. This is achieved by lifting patches of functions to an embedding dimension using a nonlocal linear operator, namely, an integral operator. Unlike ViT, the resulting neural operator architecture, which we call ViT neural operator (ViTNO), is discretization invariant and allows for zero-shot generalization to different resolutions. In Subsection 4.4 we describe a different patching-based transformer neural operator acting on function space, the Fourier attention neural operator (FANO). The architecture is discretization invariant and is based on $\mathsf{Q},\mathsf{K},\mathsf{V}$, learnable parameters in the attention mechanism, being defined as integral operators acting themselves on function space. We demonstrate in Section 6 that the approach involving parametrizing $\mathsf{Q},\mathsf{K},\mathsf{V}$ as integral operators yields an architecture with greater parameter count for a similar computational complexity. We show that this architecture produces improved results in the context of smooth problems.
@@ 4.1 Set-Up
for any $v\in\mathcal{V}$. Here the notation $\mathsf{A}_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ indicates the ambiguity in the definition of self-attention. We note that we will make explicit which definition of self-attention operator from the previous sections we employ for each architecture; we will also make clear the definitions of the parameters $\theta_{h}\in\Theta$ defining each attention operator. The operators $\mathsf{W}_{1},\mathsf{W}_{2}:\mathcal{V}\to\mathcal{V}$ are pointwise linear operators. [^4] We note that for all experiments in Section 6 we set $\mathsf{W}_{1},\mathsf{W}_{2}$ to be the identity operators. Next, the operator $\mathsf{F}_{\textrm{LayerNorm}}:\mathcal{V}\to\mathcal{V}$ defines the layer normalization and the operator $\mathsf{F}_{\textrm{NN}}:\mathcal{V}\to\mathcal{V}$ defines a feed-forward neural network layer. Both operators are applied pointwise to the input; further details are provided in the context of each architecture.
@@
Assuming ${\kappa}_{\theta}(x,y)={\kappa}_{\theta}(x-y)$, hence the kernel to be periodic, then (12) may be viewed as a convolution operator, which makes it possible to use the Fast Fourier Transform (FFT) to compute (12) and to parametrize $\mathsf{W}$ in Fourier space. Following Li et al. (2021), this insight leads to the following formulation of the operator given in Definition 19.
@@
For uniform discretizations $n_{1}\times\cdots\times n_{d}=N$ of the space $D$, the Fourier transform can be replaced with the Fast Fourier Transform (FFT). Indeed, for the Fourier transform we have $\mathcal{F}u:\mathbbm{R}^{d}\to\mathbbm{C}^{r}$ and $R_{\mathsf{W}}(\theta):\mathbbm{R}^{d}\to\mathbbm{C}^{r^{\prime}\times r}$. On the other hand, for the FFT we have $\mathcal{F}u:\mathbbm{Z}^{d}\to\mathbbm{C}^{r}$ and $R_{\mathsf{W}}(\theta):\mathbbm{Z}^{d}\to\mathbbm{C}^{r^{\prime}\times r}$. In practice, it is possible to select a finite-dimensional parametrization by choosing a finite set of wave numbers
$\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}\coloneqq\bigl{|}\bigl{\{}k\in\mathbbm{Z}^{d}:|k_{j}|\leq k_{\textrm{max},j},\,\text{for}\,j=1,\ldots,d\bigr{\}}\bigr{|}.$
We can therefore parametrize $R_{\mathsf{W}}$ by a complex $\bigl{(}(2k_{\text{max},1}+1)\times\ldots\times(2k_{\text{max},d}+1)\times r^{\prime}\times r\bigr{)}$-tensor. We will use the notation $k_{\textrm{max}}=\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}$ when discussing parameter scalings in the subsequent sections.
@@ 4.2 Transformer Neural Operator
Next, we define the details of $\mathsf{E}_{L}:\mathcal{V}\to\mathcal{V}$ in (10) by specifying a definition of $\mathsf{A}_{\textrm{MultiHead}}$, $\mathsf{W}_{1},\mathsf{W}_{2}$, $\mathsf{F}_{\textrm{LayerNorm}}$, and $\mathsf{F}_{\textrm{NN}}$ for the transformer neural operator setting. The operator $\mathsf{A}_{\textrm{MultiHead}}:\mathcal{V}\to\mathcal{V}$ is the multi-head attention operator from (11) based on the continuum self-attention from Definition 5. Letting $H\in\mathbbm{N}$ denote the number of attention heads, the multi-head attention operator is parametrized by the learnable linear transformations $Q_{h}\in\mathbbm{R}^{d_{K}\times d_{\textrm{model}}},K_{h}\in\mathbbm{R}^{d_{K}\times d_{\textrm{model}}},V_{h}\in\mathbbm{R}^{d_{K}\times d_{\textrm{model}}}$, for $h=1,\ldots,H$, where for implementation purposes $d_{K}$ is chosen as $d_{K}\coloneqq d_{\textrm{model}}/H$. [^5] We note that $d_{\textrm{model}}$ and $H$ should be chosen so that $d_{\textrm{model}}$ is a multiple of $H$. The multi-head attention operator is thus defined by the application of a linear transformation $W_{\textrm{MultiHead}}\in\mathbbm{R}^{d_{\textrm{model}}\times H\cdot d_{K}}$ to the concatenation of the outputs of $H\in\mathbbm{N}$ self-attention operators. The operators $\mathsf{W}_{1},\mathsf{W}_{2}:\mathcal{V}\to\mathcal{V}$ are pointwise linear operators and hence admit finite-dimensional representations $W_{1},W_{2}\in\mathbbm{R}^{d_{\textrm{model}}\times d_{\textrm{model}}}$, so that they define a map
@@
for $k=1,\dots,d_{\textrm{model}}$, any $v\in\mathcal{V}$, and any $x\in D\subset\mathbbm{R}^{d}$, where the notation $(\,\cdot\,)_{k}$ is used to denote the $k$’th entry of the vector. In equation (16), $\epsilon\in\mathbbm{R}^{+}$ is a fixed parameter, $\gamma_{k},\beta_{k}\in\mathbbm{R}$ are learnable parameters and $m,\sigma$ are defined as
$m(y)$ $\coloneqq\frac{1}{d_{\textrm{model}}}\sum_{k=1}^{d_{\textrm{model}}}y_{k},$ (17)
@@ 4.3 Vision Transformer Neural Operator
Let $D\subset\mathbbm{R}^{d}$ be a bounded open set. Consider $u\in\mathcal{U}(D;\mathbbm{R}^{d_{u}})$ and let $D\coloneqq D_{1}\cup\cdots\cup D_{P}$ be a uniform partition of the space $D$ so that $D_{j}\cong D^{\prime}$ have equivalent finite volume for all $j\in\llbracket 1,P\rrbracket$ for some $D^{\prime}\subset D$, where $P\in\mathbb{N}$ represents the number of patches. We consider model inputs $u\in\mathcal{U}(D;\mathbbm{R}^{d_{u}})$, model embedding space $\mathcal{V}:=\left\{v\mid v:\llbracket 1,P\rrbracket\to\mathcal{U}\bigl{(}D^{\prime};\mathbbm{R}^{d_{\textrm{model}}}\bigr{)}\right\}$, which is a space of functions acting on patch indices, and model output space $\mathcal{Z}(D;\mathbbm{R}^{d_{z}})$, with domain $D\subset\mathbbm{R}^{d}$.
@@
for $k=1,\dots,d_{\textrm{model}}$, any $p\in\llbracket 1,P\rrbracket$ and any $x\in\mathbbm{R}^{d}\subset D$, where we use the notation $(\,\cdot\,)_{k}$ to denote the $k$’th entry of the vector. In equation (24), $\epsilon\in\mathbbm{R}^{+}$ is a fixed parameter, $\gamma_{k},\beta_{k}\in\mathbbm{R}$ are learnable parameters and $m,\sigma$ are defined as in (17). The operator $\mathsf{F}_{\textrm{NN}}:\mathcal{V}\to\mathcal{V}$ is defined by
@@
Discontinuities may arise when the patching architecture is used and may be visible in error plots and can also interfere with self-composition of maps learned as solution operators of time-dependent PDEs. Ameliorating such discontinuities can be achieved using a smoothing operator as a final layer of the network. To this end, assuming $\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}\subset H^{s}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}$ for some $s\geq 0$, we define for $\epsilon>0$ and $\alpha>1$ the operator $(I-\epsilon\Delta)^{-\alpha}:\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}\to H^{s+2\alpha}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}$, and define the last layer of the architecture as
${z}\mapsto(I-\epsilon\Delta)^{-\alpha}(z),$ (29)